# Euclid Quick Data Release (Q1)

### v1

@@ 4.4 Model 4: Zoobot
Model 4 is Zoobot, which has been previously described. Refer to Sect. 3 for more details.
@@ 4.5 Model 5: LensVision
The data used for training includes S1 and S2, as well as a set of known grade A and B lenses for positives, and the random galaxies matching S2 and the labelled DESI non-lenses as the negatives. This data set was augmented to increase the size by including rotations and corner cutouts of each image. The images used the scaling on the -only images.
@@ 5 Results
First, a group of strong lensing experts within the Euclid Collaboration visually inspected the top ranked objects from each of the models 1–5, in what is referred to as the Galaxy Judges (GJ) project (Euclid Collaboration: Walmsley et al. 2025). This was to gain an initial understanding of how the models compared in lens finding, and was to inform which models should be prioritised in selecting which candidates were shown to citizens. Following this was the citizen science visual inspection, for which we created a Euclid-specific version of the Space Warps (SW) project ran on the Zooniverse platform (Marshall et al. 2015; More et al. 2016). All objects that were ranked highly by these citizens were then added to the GJ project and were ranked by experts, giving the final grades for all the interesting lens-like objects discovered. In both the SW citizen science project and the GJ expert classifications, each image was inspected by multiple humans, with the results being aggregated in order to reduce noise and the level of subjectivity. For SW, the citizens were given the option of lens or not a lens, while in GJ the experts assigned a grade to quantify the quality of a lens. These grades were: X (not a lens); C (contains lens-like features but not obviously a lens); B (probably a lens); and A (confidently a lens). Examples of lenses with these grades are shown in Appendix B. In this paper we focus on using the results from the visual inspection to quantify model performance; for a detailed analysis of the visual inspection process see Euclid Collaboration: Walmsley et al. (2025).
@@ 5.1 Preliminary expert classifications for model comparison
Each of the five models’ top objects were ranked by experts, with each cutout receiving approximately 10 votes which were then aggregated into one grade. Note that a rank of corresponds to different threshold scores for each machine-learning model: the corresponding scores for Models 1–5 are 0.992, 0.999, 0.983, 0.999, and 0.967, respectively. The aim of this was to evaluate each model’s ability to find lenses. The two main areas of consideration in this analysis were: (1) how many lenses of each grade were found by each model; and (2) of the lenses that one model did not rank in its top object but were highly ranked by another model, where did these missed lenses sit in its score distribution. The first of these is addressed in Fig. 5, displaying the number of grade A and the number of grade A + B lenses found by each of the Models 1–5. The breakdown of the number of lenses and non-lenses in each model’s top 1000 ranked objects is shown in Table 2. The best performing model is Model 4 (Zoobot), discovering almost twice the number of grade A lenses as the second-best performing model, model 1 (4-Layer CNN). Although the other models discover fewer lenses, the proportion of lenses picked out by every model in the top 1000 is significantly higher than the random lens to non-lens ratio (around 1:). However, the top 1000 ranked images from every model are still non-lenses by majority, meaning visual inspection by humans is essential.
@@ 5.2 Citizen scientist classifications
Once we had verified that all the models were able to find lenses, we picked the images that were ranked highly by any of the models and uploaded these to a public citizen science SW project on the Zooniverse platform. Here, citizens classified each cutout based on whether they thought it contained a lens or not. This took place over the span of about six weeks, with total classifications being made by approximately users. Images that were flagged as possible lenses by the citizens were then passed to expert inspection in the GJ project to get a final grade. It was decided to use SW to visually inspect the top objects from each model, and once this was completed the next from Model 1 and Model 4 were added, along with a subset of the next images from the other models.444The specifics of SW are outlined in the system overview paper (Euclid Collaboration: Walmsley et al. 2025). A random sample of images was also added for calibration.
Of the cutouts that we started with, a total of objects were inspected by SW, of which a subset of were graded highly enough to make it to the follow-up expert classification stage. Of these, there were 250 grade A, 247 grade B, 585 grade C lenses, and 6280 non-lenses. The top grade A lenses are displayed in Sect. A.1. Of these, only four grade B and five grade C Q1 lenses appeared in the ‘known real lenses’ data set that some models used for training. Almost all the lenses discovered were from images that were uploaded to SW due to being highly ranked by at least one of the machine-learning models: grade A lenses and grade B lenses were found in the random images that were not selected for SW by any of the machine-learning models. A distribution of how all these lenses were ranked by each model is shown in Fig. 6. The catalogue of all the lenses we found is publicly available on Zenodo555https://doi.org/10.5281/zenodo.15003116.
@@ 5.3 Completeness of our lens sample
In addition to uploading highly ranked images that had been selected by the machine-learning models, we also uploaded randomly-selected images from the original Q1 images. This allowed us to quantify how good the machine-learning models are at finding lenses and forecast how many lenses we would have found had we visually inspected the full Q1 sample. Note that this random sample necessarily had a small overlap with the machine-learning-selected sample, since otherwise it would have contained less lenses than average. Within this random sample we found 9 grade A lenses and 22 grade B lenses, although many of these random images had also been selected by the machine-learning models.
We can use Bayes’ theorem to infer the rate of lenses in the total sample, given that we found 9 grade A and 22 grade B lenses in a sample of . Assuming a log-uniform prior on the rate of lenses between 0 and 1 we infer grade A lenses and grade B lenses (68% confidence). The overall system found 250 grade A lenses and 247 grade B lenses (497 grade A + B), indicating that some grade B lenses may have been missed, but that it is likely that we discovered almost all grade A lenses. A substantially better estimate of the total number of lenses in Q1 can be made by including the fact that we have complete knowledge of how the system performs up to a Zoobot rank of . This is the regime where most lenses exist, but it is sparsely probed by the random sample. Given that the random sample included 2 grade A lenses beyond this point, we can infer that Zoobot missed grade A lenses (68% confidence). For grade B, Zoobot’s top misses lenses. The combination of the other four machine-learning models found 26 of the grade A lenses and 68 of the grade B lenses that Zoobot missed.
The binomial estimates in the previous paragraph allow us to estimate where the Zoobot ROC curve should end. We do not know how the ROC curve should evolve between and the final image, but it is likely that the rate of missing grade A and B lenses decreases with Zoobot rank. Figure 7 shows the range of ROC-like curves for Zoobot that are consistent with the binomial estimates and assuming the probability of finding a lens is proportional to the Zoobot rank between and the final image.
Furthermore, it is important to understand the completeness of this sample as a function of properties such as brightness and Einstein radius. An analysis of this for the full Q1 lens catalogue is presented in Euclid Collaboration: Walmsley et al. (2025), but we will be able to do this for specific machine learning models in the future when we have enough real lenses to sufficiently span this parameter space.
@@ 6.1 Machine-learning takeaways
A similar case is that of image scalings, which again varied across models and did not appear to directly correlate with machine-learning performance. Zoobot found that when using -only images, the scaling resulted in a better performance than MTF, while Model 2 found that the scaling picked out more stars and performed worse than the MTF scaling. Models 3 and 5 used the -only images, and Model 1 trained -only models for both the and MTF scaling and averaged the results. Although colour images were available, all teams decided to use the -only greyscale images in the best versions of their models, despite many teams also testing colour-based models. This is of interest as during human visual inspection, it is often found that colour aids in the detection of lenses, since there is typically a red/blue gradient between the lens/source galaxy that is helpful for distinguishing lenses. The fact that no team found there to be a significant improvement in performance when adding colour data may suggest that machine-learning models prioritise different aspects of images when classifying lenses, or alternatively that these models are not able to pick up on properties such as colour gradients as well as they are able to pick up on morphological information. The latter may make sense in the case of Zoobot; through being pretrained on galaxies of different morphologies, it is not clear that there would be any pattern in colour information that could be extrapolated to lenses in the way that patterns picked up in morphology may be useful. Patterns picked up in the difference between intensities in different filter bands likely might be something that happens in the first few layers of a network and therefore is not something that fine-tuning can address.
Figure 8 shows the position of grade A lenses as a function of Model 1 and Model 4 rank. The two ranks are broadly correlated, but with substantial scatter. It is not clear if the scatter is due to the networks learning different features, or due to the intrinsic uncertainty of each model. We colour each of the points according to the model that best ranked each grade A lens. This can be seen as a marker for which network would have discovered it first. Models 1 and 4 combined make up the large majority of the grade A lens first discoveries (88) and the combination of them results in a powerful lens finder. Many grade A lenses beyond Model 1’s top were highly ranked by Model 4 and vice versa. While Models 2, 3 and 5 were able to provide additional lenses, only 7 were beyond the inspection threshold of Models 1 and 4. These results suggest that Models 1 and 4 provide the bulk of the discovery information. Future searches should use a verification set to check machine-learning performance and use that performance to prioritise the visual inspection effort between networks, or to build an ensemble classifier (Euclid Collaboration: Holloway et al. 2025).
@@ 6.2 The future of lens finding in Euclid
With the number of lenses discovered thus far, it is evident that a large number of strong lenses in Euclid are discoverable, and the future releases of Euclid data are going to contain a wealth of interesting data for strong lensing. However, while the number of lenses will scale with the amount of data Euclid collects, the number of objects that can be visually inspected by humans will remain roughly constant, providing a challenge for lens finding as long as humans remain better at recognising strong gravitational lenses than machine-learning algorithms.
The number of objects able to be classified by humans in the SW project was around , meaning 10% of the Q1 data was visually inspected. However, the Euclid DR1 data will cover around deg2 of the EWS ( times the area of Q1), and hence visually inspecting objects will only make up 0.25% of the total DR1 data. Previous lens-finding citizen-science projects have been able to search through a larger number of images (e.g., in the CFHT Legacy Survey SW project; see More et al. 2016), so assuming an upper limit of visually inspecting one million objects, it is therefore informative to investigate what lenses fall inside/outside the top 0.2%–2% of Q1 (corresponding to the top – objects).
In the left panel of Fig. 9 we present the density of lenses found in Q1 as a function of Zoobot rank. Although there will be lenses in Q1 that we did not discover, as explained in Sect. 5.3 almost all grade A lenses were likely discovered and hence we can take the frequency of the discovered grade A lenses as being the true frequency. The worst rank of a grade A lens according to Zoobot is , meaning Zoobot alone is unable to discover 100% of the grade A lenses in DR1. However, 152 (220) of the grade A lenses were ranked in the top 2000 () by Zoobot, meaning the number of grade A lenses in DR1 that we will be able to discover with the help of citizen science is around –, which will be by far the largest collection of strong lens candidates at the time. Looking beyond DR1, in the full Euclid 6-year survey the top million ranked Zoobot objects will contain grade A/B lenses, providing a revolution for strong lensing.
@@ 6.3 Discovery of lenses of particular interest
An ensemble approach to lens finding will still benefit from the individual machine-learning models being optimised as much as possible, and one aspect that is worth considering is the ability to discover lenses of individual scientific value. Figure 9 shows the rank and the corresponding images of objects that may be more interesting than the typical single-galaxy-lensed-by-an-LRG lensing configuration that dominate the strong lens samples and are typically the only type of lens system that is simulated. It is interesting to investigate how these objects are scored by machine-learning models which are not trained to recognise lenses of this type.
@@ 6.3.1 DSPLs
DSPLs are useful cosmological probes; an analysis and modelling of the DSPLs found in Euclid Q1 is presented by Euclid Collaboration: Li et al. (2025). There it was found that four lenses in Q1 could be convincingly modelled as DSPLs, and there were another four candidate DSPLs. The ranking of the four DSPLs by each of the models is presented in Table 3, and images of these four DSPLs are shown in Sect. 11. Although four DSPLs is not enough to confidently extrapolate about our ability to discover DSPLs in the EWS, it is reassuring that many of them were ranked highly. Model 1 performs exceedingly well at picking out the DSPLs, with all four of them being ranked in the top 500. With the combined performance of the five models it is unlikely that we missed many high-quality DSPLs, with only forecasted to be detectable within Q1 to begin with (Euclid Collaboration: Li et al. 2025).
Although worse than Model 1, Zoobot also performs well at recovering DSPLs, with two of the four DSPLs being in the top 10 ranked objects (), at ranks 3 and 7. The ranks of the other two DSPLs were 1902 and 2349, which although sitting at the lower end of the grade A lens ranks, still ranked highly enough for all of them to likely be recovered if they were in DR1. Of the four DSPLs, the two that were ranked in the top 10 both have a larger Einstein radii, with brighter arcs that are more easily distinguishable than the two lower ranked. The DSPL candidates on average rank lower, with ranks of 622, 1540, 4277, and , and also have smaller Einstein radii, fainter arcs and/or lens light blending with the arc light. Although only candidate DSPLs, these are still systems that we are interested in, and the fact that these are all ranked on the lower end amongst the lenses ranked by Zoobot (though still in the top 2.5% of the whole sample), this may be something we want to address in future machine-learning training.

### v2

@@ 4.4 Model 4: Zoobot
Model 4 is Zoobot, which has been described previously. (Refer to Sect. 3 for more details.)
@@ 4.5 Model 5: LensVision
The data used for training includes S1 and S2 as well as a set of known grade A and B lenses for positives and the random galaxies matching S2 and the labelled DESI non-lenses as the negatives. This data set was augmented to increase the size by including rotations, horizontal and vertical flips, and corner cutouts cropping the original images to . These augmentations resulted in images in both the positive and negative class for training. The images used the scaling on the -only images.
@@ 5 Results
First, a group of strong lensing experts within the Euclid Collaboration visually inspected the top ranked objects from each of the models 1–5, in what is referred to as the Galaxy Judges (GJ) project (Euclid Collaboration: Walmsley et al. 2025a). This was to gain an initial understanding of how the models compared in lens finding, and was to inform which models should be prioritised in selecting which candidates were shown to citizens. Following this was the citizen science visual inspection, for which we created a Euclid-specific version of the Space Warps (SW) project ran on the Zooniverse platform (Marshall et al. 2015; More et al. 2016). All objects that were ranked highly by these citizens were then added to the GJ project and were ranked by experts, giving the final grades for all the interesting lens-like objects discovered. In both the SW citizen science project and the GJ expert classifications, each image was inspected by multiple humans, with the results being aggregated in order to reduce noise and the level of subjectivity. For SW, the citizens were given the option of lens or not a lens, while in GJ the experts assigned a grade to quantify the quality of a lens. These grades were: X (not a lens); C (contains lens-like features but not obviously a lens); B (probably a lens); and A (confidently a lens). Examples of lenses with these grades are shown in Appendix B. In this paper we focus on using the results from the visual inspection to quantify model performance; for a detailed analysis of the visual inspection process see Euclid Collaboration: Walmsley et al. (2025a).
@@ 5.1 Preliminary expert classifications for model comparison
Each of the five models’ top objects were ranked by experts, with each cutout receiving approximately ten votes which were then aggregated into one grade. We note that a rank of corresponds to different threshold scores for each machine-learning model: the corresponding scores for Models 1–5 are 0.992, 0.999, 0.983, 0.999, and 0.967, respectively. The aim of this was to evaluate each model’s ability to find lenses. The two main areas of consideration in this analysis were: (1) how many lenses of each grade were found by each model; and (2) of the lenses that one model did not rank in its top object but were highly ranked by another model, where did these missed lenses sit in its score distribution. The first of these is addressed in Fig. 8, displaying the number of grade A and the number of grade A + B lenses found by each of the Models 1–5. The breakdown of the number of lenses and non-lenses in each model’s top 1000 ranked objects is shown in Table 2. The best performing model is Model 4 (Zoobot), discovering almost twice the number of grade A lenses as the second-best performing model, Model 1 (four-layer CNN). Although the other models discover fewer lenses, the proportion of lenses picked out by every model in the top 1000 is significantly higher than the random lens to non-lens ratio (around 1:). However, the top 1000 ranked images from every model are still non-lenses by majority, meaning visual inspection by humans is essential.
@@ 5.2 Citizen scientist classifications
Once we had verified that all the models were able to find lenses, we picked the images that were ranked highly by any of the models and uploaded these to a public citizen science SW project on the Zooniverse platform. Here, citizens classified each cutout based on whether they thought it contained a lens or not. This took place over the span of about six weeks, with total classifications being made by approximately users. Images that were flagged as possible lenses by the citizens were then passed to expert inspection in the GJ project to get a final grade. It was decided to use SW to visually inspect the top objects from each model, and once this was completed the next from Model 1 and Model 4 were added, along with a subset of the next images from the other models.555The specifics of SW are outlined in the system overview paper (Euclid Collaboration: Walmsley et al. 2025a). A random sample of images was also added for calibration.
Of the cutouts that we started with, a total of objects were inspected by SW, of which a subset of were graded highly enough to make it to the follow-up expert classification stage. Of these, there were 250 grade A, 247 grade B, 585 grade C lenses, and 6280 non-lenses. The top grade A lenses are displayed in Sect. A.1. Of these, only four grade B and five grade C Q1 lenses appeared in the ‘known real lenses’ data set that some models used for training. Almost all the lenses discovered were from images that were uploaded to SW due to being highly ranked by at least one of the machine-learning models: grade A lenses and grade B lenses were found in the random images that were not selected for SW by any of the machine-learning models. A distribution of how all these lenses were ranked by each model is shown in Fig. 9. The catalogue of all the lenses we found is publicly available on Zenodo666https://doi.org/10.5281/zenodo.15003116.
@@ 5.3 Completeness of our lens sample
In addition to uploading highly ranked images that had been selected by the machine-learning models, we also uploaded randomly selected images from the original Q1 images. This allowed us to quantify how good the machine-learning models are at finding lenses and forecast how many lenses we would have found had we visually inspected the full Q1 sample. We note that this random sample necessarily had a small overlap with the machine-learning-selected sample, since otherwise it would have contained fewer lenses than average. Within this random sample we found nine grade A lenses and 22 grade B lenses, although many of these random images had also been selected by the machine-learning models.
We used Bayes’ theorem to infer the rate of lenses in the total sample, given that we found nine grade A and 22 grade B lenses in a sample of . Assuming a log-uniform prior on the rate of lenses between 0 and 1, we inferred grade A lenses and grade B lenses (68% confidence). The overall system found 250 grade A lenses and 247 grade B lenses (497 grade A + B), indicating that some grade B lenses may have been missed, but it is likely that we discovered almost all grade A lenses. A substantially better estimate of the total number of lenses in Q1 can be made by including the fact that we have complete knowledge of how the system performs up to a Zoobot rank of . This is the regime where most lenses exist, but it is sparsely probed by the random sample. Given that the random sample included two grade A lenses beyond this point, we can infer that Zoobot missed grade A lenses (68% confidence). For grade B, Zoobot’s top misses lenses. The combination of the other four machine-learning models found 26 of the grade A lenses and 68 of the grade B lenses that Zoobot missed.
The binomial estimates in the previous paragraph allowed us to estimate where the Zoobot ROC curve should end. We do not know how the ROC curve should evolve between and the final image, but it is likely that the rate of missing grade A and B lenses decreases with Zoobot rank. Figure 10 shows the range of ROC-like curves for Zoobot that are consistent with the binomial estimates and assuming the probability of finding a lens is proportional to the Zoobot rank between and the final image.
Furthermore, it is important to understand the completeness of this sample as a function of properties such as brightness and Einstein radius. An analysis of this for the full Q1 lens catalogue is presented in Euclid Collaboration: Walmsley et al. (2025a), but we will be able to do this for specific machine learning models in the future when we have enough real lenses to sufficiently span this parameter space.
@@ 6.1 Machine-learning takeaways
A similar case is that of image scalings, which again varied across models and did not appear to directly correlate with machine-learning performance. Zoobot found that when using -only images, the scaling resulted in a better performance than MTF, while Model 2 found that the scaling picked out more stars and performed worse than the MTF scaling. Models 3 and 5 used the -only images, and Model 1 trained -only models for both the and MTF scaling and averaged the results. Although colour images were available, all teams decided to use the -only greyscale images in the best versions of their models, despite many teams also testing colour-based models. This is of interest as during human visual inspection, it is often found that colour aids in the detection of lenses, since there is typically a red–blue gradient between the lens and source galaxy that is helpful for distinguishing lenses. The fact that no team found there to be a significant improvement in performance when adding colour data may suggest that machine-learning models prioritise different aspects of images when classifying lenses, or alternatively that these models are not able to pick up on properties such as colour gradients as well as they are able to pick up on morphological information. The latter may make sense in the case of Zoobot; through being pretrained on galaxies of different morphologies, it is not clear that there would be any pattern in colour information that could be extrapolated to lenses in the way that patterns picked up in morphology may be useful. Patterns picked up in the difference between intensities in different filter bands likely might be something that happens in the first few layers of a network and therefore is not something that fine-tuning can address.
Figure 11 shows the position of grade A lenses as a function of Model 1 and Model 4 rank. The two ranks are broadly correlated, but with substantial scatter. It is not clear if the scatter is due to the networks learning different features, or due to the intrinsic uncertainty of each model. We colour each of the points according to the model that best ranked each grade A lens. This can be seen as a marker for which network would have discovered it first. Models 1 and 4 combined make up the large majority of the grade A lens first discoveries (88) and the combination of them results in a powerful lens finder. Many grade A lenses beyond Model 1’s top were highly ranked by Model 4 and vice versa. While Models 2, 3, and 5 were able to provide additional lenses, only seven were beyond the inspection threshold of Models 1 and 4. These results suggest that Models 1 and 4 provide the bulk of the discovery information. Future searches should use a verification set to check machine-learning performance and use that performance to prioritise the visual inspection effort between networks, or to build an ensemble classifier (Euclid Collaboration: Holloway et al. 2025).
@@ 6.2 The future of lens finding in Euclid
With the number of lenses discovered thus far, it is evident that a large number of strong lenses in Euclid are discoverable, and the future releases of Euclid data are going to contain a wealth of interesting data for strong lensing. However, while the number of lenses will scale with the amount of data Euclid collects, the number of objects that can be visually inspected by humans will remain roughly constant, providing a challenge for lens finding as long as humans remain better at recognising strong gravitational lenses than machine-learning algorithms. Compared to visual inspection, the machine learning computation time is negligible, taking Zoobot less than an hour to run on Q1 using an NVIDIA A100 80GB GPU.
The number of objects able to be classified by humans in the SW project was around , meaning 10% of the Q1 data was visually inspected. However, the Euclid DR1 data will cover around deg2 of the EWS ( times the area of Q1), and hence visually inspecting objects will only make up 0.25% of the total DR1 data. Previous lens-finding citizen-science projects have been able to search through a larger number of images (e.g. in the CFHT Legacy Survey SW project; see More et al. 2016), so assuming an upper limit of visually inspecting one million objects, it is therefore informative to investigate what lenses fall inside and outside the top 0.2%–2% of Q1 (corresponding to the top – objects).
In the left panel of Fig. 12 we present the density of lenses found in Q1 as a function of Zoobot rank. Although there will be lenses in Q1 that we did not discover, as explained in Sect. 5.3, almost all grade A lenses were likely discovered, and hence we can take the frequency of the discovered grade A lenses as being the true frequency. The worst rank of a grade A lens according to Zoobot is , meaning Zoobot alone is unable to discover 100% of the grade A lenses in DR1. However, 152 (220) of the grade A lenses were ranked in the top 2000 () by Zoobot, meaning the number of grade A lenses in DR1 that we will be able to discover with the help of citizen science is around –, which will be by far the largest collection of strong lens candidates at the time. Looking beyond DR1, in the full Euclid 6-year survey the top million ranked Zoobot objects will contain grade A+B lenses, providing a revolution for strong lensing.
@@ 6.3 Discovery of lenses of particular interest
An ensemble approach to lens finding will still benefit from the individual machine-learning models being optimised as much as possible, and one aspect that is worth considering is the ability to discover lenses of individual scientific value. Figure 12 shows the rank and the corresponding images of objects that may be more interesting than the typical single-galaxy-lensed-by-an-LRG lensing configuration that dominate the strong lens samples and are typically the only type of lens system that is simulated. It is interesting to investigate how these objects are scored by machine-learning models which are not trained to recognise lenses of this type, such as DSPLs and non-LRG lens galaxies. We note that one population of strong lenses that are underrepresented in our sample are lensed quasars, and hence a designated search is required to discover the – lensed quasars expected in the full EWS.
@@ 6.3.1 DSPLs
DSPLs are useful cosmological probes; an analysis and modelling of the DSPLs found in Euclid Q1 is presented by Euclid Collaboration: Li et al. (2025). There it was found that four lenses in Q1 could be convincingly modelled as DSPLs, and there were another four candidate DSPLs. The ranking of the four DSPLs by each of the models is presented in Table 3, and images of these four DSPLs are shown in Sect. 14. Although four DSPLs is not enough to confidently extrapolate about our ability to discover DSPLs in the EWS, it is reassuring that many of them were ranked highly. Model 1 performs exceedingly well at picking out the DSPLs, with all four of them being ranked in the top 500. With the combined performance of the five models it is unlikely that we missed many high-quality DSPLs, with only forecasted to be detectable within Q1 to begin with (Euclid Collaboration: Li et al. 2025).
Although worse than Model 1, Zoobot also performs well at recovering DSPLs, with two of the four DSPLs being in the top 10 ranked objects (), at ranks 3 and 7. The ranks of the other two DSPLs were 1902 and 2349, which although sitting at the lower end of the grade A lens ranks, still ranked highly enough for all of them to likely be recovered if they were in DR1. Of the four DSPLs, the two that were ranked in the top 10 both have a larger Einstein radii, with brighter arcs that are more easily distinguishable than the two lower ranked. The DSPL candidates on average rank lower, with ranks of 622, 1540, 4277, and , and also have smaller Einstein radii, fainter arcs or lens light blending with the arc light. Although only candidate DSPLs, these are still systems that we are interested in, and the fact that these are all ranked on the lower end amongst the lenses ranked by Zoobot (though still in the top 2.5% of the whole sample), this may be something we want to address in future machine-learning training.